OpenStudy (anonymous):

Evaluate. \[\lim_{x \rightarrow \infty}\left[ \log_{5}(\frac{ 1 }{ 125 }-2^{-x}) \right]\]

OpenStudy (rational):

log is continuous function, so you can send the limit inside the function : \[\lim(\log(f(x))) = \log(\lim f(x))\]

OpenStudy (rational):

\[\lim_{x \rightarrow \infty}\left[ \log_{5}(\frac{ 1 }{ 125 }-2^{-x}) \right] = \log_5\left[ \color{blue}{\lim_{x \rightarrow \infty}(\frac{ 1 }{ 125 }-2^{-x})} \right]\]

OpenStudy (anonymous):

oh, okay

OpenStudy (anonymous):

What do I do next?

OpenStudy (rational):

you can do many things maybe just think of what happens to \(\large 2^{-x}\) as you make \(x\) large

OpenStudy (rational):

\(2^{-1} = ?\) \(2^{-2} = ?\) \(2^{-3} = ?\) \(\cdots\) \(2^{-100} = ?\)

OpenStudy (anonymous):

Doesn't the value become smaller?

OpenStudy (rational):

evaluate those values and see

OpenStudy (anonymous):

.5 .25 .125 7.8888...E-31

OpenStudy (rational):

you can see the value of \(2^{-x}\) is approaching \(0\) as you increase \(x\) so \[\lim\limits_{x\to\infty}2^{-x} = 0\]

OpenStudy (rational):

\[\begin{align} \lim_{x \rightarrow \infty}\left[ \log_{5}(\frac{ 1 }{ 125 }-2^{-x}) \right] &= \log_5\left[ \color{blue}{\lim_{x \rightarrow \infty}(\frac{ 1 }{ 125 }-2^{-x})} \right]\\~\\ &=\log_5\left[ \color{blue}{\frac{ 1 }{ 125 }-0} \right]\\~\\ &=\log_5\left[ \color{blue}{5^{-3}} \right]\\~\\ &=-3 \end{align}\]

OpenStudy (anonymous):

oohh! I see. Thanks for explaining this. Greatly appreciated! :) I'm not really good at limits and thinking about infinities and such. Any tips?

OpenStudy (rational):

my only tip is not to try and visualize everything, sometimes you need to just follow the rules and things will be simple

OpenStudy (rational):

not meant to say, stop visualizing... just want to say that following rules is also important as calculus is very huge, learning wont be smooth w/o a systematic approach graph everything but don't always try to understand in terms of graphs only https://www.desmos.com/calculator

OpenStudy (anonymous):

Oh, okay. I will keep that in mind. Thanks for everyting! :)